OpenStudy (anonymous):

I am having trouble with this simultaneous equation y = x^2 + 1 2y + 2 = x^2 + 6x

OpenStudy (anonymous):

\[y=x^2+1\] or: \[x^2=y-1\] plug in this value of x^2 into the second equation to get:\[2y+2=(y-1)+6x\]

OpenStudy (anonymous):

which would give you \[y=6x-3\]

OpenStudy (anonymous):

which you can then plug into either equation and solve for x using quadratic formula